OpenStudy (anonymous):

Solve the following system of equations. 1x+2y=5 -3x-6y=-15

OpenStudy (anonymous):

I get 0=0. Have you learned matrices yet?

OpenStudy (mathlegend):

@RoLexx do you want me to show the work?

OpenStudy (anonymous):

I got the same thing however how when i input the answer in its telling me im incorrect

OpenStudy (mathlegend):

0=0 is No Solution

OpenStudy (anonymous):

0=0 actually means infinite amount of solutions

OpenStudy (mathlegend):

It is No Solution.

OpenStudy (mathlegend):

I just looked in my textbook just to make sure before I posted.

OpenStudy (anonymous):

The directions say that If there are no solutions, type "No Solution" for both and . If there are infinitely many solutions, type "x" forx , and an expression in terms of x for y.

OpenStudy (anonymous):

I've tried several times and I input x for x which is correct but what expression do i need to input in terms of x for y?

OpenStudy (mathlegend):

Maybe my book is wrong.

OpenStudy (anonymous):

x + 2y = 5 -3x - 6y = -15 Setting up a matrix 1 2 | 5 -3 -6 | -15 add (3 times row1) to row 2 1 2 | 5 0 0 | 0 y is arbitrary. x + 2y = 5 y = s x = -2s + 5 y = s True for all values of s. Infinite values of s answer in matrix form | x | = s | -2 | + 5 | 1 | | y | | 1 | | 0 |

OpenStudy (mathlegend):

I actually multiplied one equation by another. I did not use a matrix. I do believe it to be No Solution.

OpenStudy (mathlegend):

Since the way I did it was specifically one of the ways you would normally do a problem like this.

OpenStudy (anonymous):

substitution x+2y=5 -3x-6y=-15 -3(-2y + 5) - 6y = -15 6y - 15 - 6y = -15 6y - 15 = 6y - 15 6y = 6y y = y for any value of y this will be correct. x will change with x of course, so infinite solutions

OpenStudy (anonymous):

you r going to have to get a third opinion @RoLexx

OpenStudy (anonymous):

I've tried that solution as well @ChmE

OpenStudy (anonymous):

But it's incorrect i thank you and @MathLegend though for the advice.

OpenStudy (anonymous):

if you put any value of y in the first equation and solve for x. you get the same x in the second equation using the same y. Infinite solutions

OpenStudy (mathlegend):

I said I could be wrong... but I looked at a problem in my textbook and it is just like this one where you get 0 = 0 and the answer says No Solution.

OpenStudy (mathlegend):

@Chme I looked in another place in my textbook and it says 0=0 is infinite solutions. So I believe @ChmE

OpenStudy (mathlegend):

Must have been read wrong by the publisher.

OpenStudy (anonymous):

So is the book saying it is no solution because it is the same line and therefor you only have one equation with 2 unknowns and it can't be solved. I'd buy that

OpenStudy (anonymous):

Just take the first one and say: If x=x then: x+2y=5 implies y=(1/2)(5-x)

OpenStudy (anonymous):

It depends how you interpret the problem

OpenStudy (anonymous):

But there is a solution, an infinite number of them. Given by the line.

OpenStudy (mathlegend):

It says that 0=0 is always true, there are infinitely many solutions.

OpenStudy (anonymous):

I would put the answer as infinite

OpenStudy (anonymous):

@malevolence19 thanks that correct but how did you get that (I'm curious)

OpenStudy (anonymous):

Technically the answer is: \[(x,y) \in \left\{(x,y) | y=\frac{1}{2}(5-x); x \in \mathbb{R}\right\}\]

OpenStudy (anonymous):

I just took the first equation and solved for y. You have x, our "parameter" so you want y in terms of that x. Since both lines are the same you can solve using either, I just used the one without the extra -3 factor.

OpenStudy (anonymous):

That explains it much better. Thanks :)